OpenStudy (anonymous):

Some clarification would be nice. Find the center and radius:

OpenStudy (anonymous):

\[2x^2-4x+2y^2-16y-12=0\]

OpenStudy (anonymous):

I got pretty far in the question but I am stuck near the end.

OpenStudy (anonymous):

mother....wow..

OpenStudy (anonymous):

It's right after I get to: \[x^2-2x+y^2-8y=6\]

OpenStudy (anonymous):

Dude, I don't even know what I'm looking at.

OpenStudy (anonymous):

center=(-g,-f) when a,b=1 and radius=(g^2+f^2-c)^.5

OpenStudy (across):

\[2x^2-4x+2y^2-16y-12=0\]\[2(x^2-2x)+2(y^2-8y)=12\]\[2(x^2-2x+1)+2(y^2-8y+16)=12+2+16\]\[(x-1)^2+(y-4)^2=15\]Does that help?

OpenStudy (across):

I meant\[(x-1)^2+(y-4)^2=23\]

OpenStudy (anonymous):

*mind blown* THANKS!

OpenStudy (anonymous):

Whoop, I see what I did wrong, I forgot to cancel the variable on the 2x. Much appreciated!